# Identifying Proca-star mergers via consistent ultralight-boson mass estimates across gravitational-wave events

### v1

@@ II.3 A boson-counting experiment:
We now generalise the above formalism to $n$ Proca-star families characterised by bosons with respective masses $\mu_{B}^{i}$ with $i\in[1,n]$. In this case, the likelihood for the boson masses and the corresponding fractions and masses $\zeta_{i}$ is given by
@@
$=\prod_{i=1}^{i=N}\left[\sum_{j=1}^{N_{\rm bosons}}p(d_{i}|\text{PSM}(\mu_{B}^{j}))\zeta_{j}+p(d_{i}|\text{BBH})\quantity(1-\sum_{j=1}^{N_{\rm bosons}}\zeta_{j})\right]$ $\propto\prod_{i=1}^{i=N}\quantity[\sum_{j=1}^{N_{\rm bosons}}{\cal{B}}^{\text{PSM}}_{\text{BBH}}(\mu_{B}^{j})\zeta_{j}+\quantity(1-\sum_{j=1}^{N_{\rm bosons}}\zeta_{j})].$
@@ V D V D V D V D
Finally, the Bayesian evidence for a model described by a number of bosons $N_{\rm bosons}=n$ is just given by
$p(\{d_{i}\}|n)=\int p(\{d_{i}\}|\{\zeta_{j},\mu_{B}^{j}\})\prod_{j=1}^{n}p(\zeta_{j},\mu_{B}^{j})\,d\zeta_{j}d\mu_{B}^{i}\ ,$ (6)
where $i\in\{1,N_{\rm bosons}=n\}$.
Finally, we can construct the posterior distribution on the number of bosons $n$ as
@@
where we will always consider an uniform prior $\pi(n)$ on $n$.
@@ II.4 Parameter sampling
We impose uniform priors on the boson masses with $\mu_{B}^{i}\in[2,15]\times 10^{-13}$ eV, imposing $\mu_{i+1}\leq\mu_{i}$. Importantly, for computational-cost reasons, we do not sample over a continuous range of values $\mu^{i}_{B}$. Instead over a discrete grid with step $\delta\mu_{B}=1\times 10^{-15}$ eV. Physically, $\delta\mu_{B}$ can be interpreted as the maximum resolution in the boson mass we allow for, which is much smaller than the standard deviation of the boson-mass posteriors from the individual that we take as input. We also set a flat prior on the respective fractions $\zeta_{i}\in[0,1]$, imposing $\sum_{i}\zeta_{i}\leq 1$. We sample the parameter space using the nested sampler Dynesty with 4096 live points.
@@ II.5 Inputs for our analysis
We take as inputs the posterior parameter distributions and Bayesian evidences for the BBH and PSM models reported in Calderón Bustillo et al. (2023), which we respectively show in Fig. 1 and Table I. In this study, the authors performed Bayesian parameter inference and model selection on the events we consider, both under the BBH and PSM scenarios. For the BBH case, the data was analyzed with the state-of-the art waveform model for BBHs NRSur7dq4 Varma et al. (2019). This model includes waveform multipoles up to $\ell=4$ and is directly calibrated to numerical simulations of precessing BBHs Boyle et al. (2019) with mass ratios $q\in[1,4]$ and spin-magnitudes $a_{i}\in[0,0.8]$ but can be extrapolated to $q=6$ and extremal spins. For the PSM case, the authors employed a family of 769 numerical simulations of unequal-mass head-on mergers of Proca stars, with the corresponding field frequencies ranging in $\omega/\mu_{B}\in[0.80,0.93]$ and imposing a flat prior on these parameters. In these configurations, the stars were released at rest at a separation distance of $20\,M$, expressed in geometric units.
@@ III Results on real data
Figure 2: Posterior probability for the number $N_{\rm bosons}$ of ultra-light bosons given Advanced LIGO - Virgo events. Solid (dashed) lines ignore (include) the event S200114. Blue and green lines respectively make use of distance priors uniform in co-moving volume and on luminosity distance.
@@
Figure 3: Posterior distributions for the boson-mass and Proca-star fraction for the single boson model. Blue and orange curves and contours denote results including excluding and including the S200114 trigger. Results are obtained assuming a distance prior uniform in co-moving volume. The posterior distribution for the boson mass has non-zero support for all the prior range because the posterior distribution for $\zeta$ is non-null at $\zeta=0$, in which case all values of $\mu_{B}$ are allowed.
@@
Figure 4: Posterior distributions for boson-mass (right) and Proca-star fractions (left) for the 2-boson model. Solid and dashed curves and contours denote results excluding and including the S200114 trigger. Blue and orange curves respectively denote the value of the primary and secondary boson-mass $\mu_{B}^{1,2}$ and corresponding fraction $\zeta_{1,2}$. Black curves denote the posterior distribution for $\zeta_{1}+\zeta_{2}$.
@@
Fig. 2 shows the posterior probability for the number of ultralight bosons $n$ given our different data sets and priors. Blue and green curves respectively exclude and include the trigger S200114. We start by discussing results under the regular distance prior. In this situation, $N_{\rm bosons}=0$ model is preferred by a factor ${\cal{B}}^{n=0}_{n=1}\simeq 1.6$ with respect to the $N_{\rm bosons}=1$ model. This weak preference $N=0$ is expected, as Table 1 shows that only one of the analyzed events has a slight preference for the PSM scenario while the other two reject such hypothesis rather strongly. In other words, the slight preference of GW190521 for the PSM hypothesis is consistent with a statistical fluctuation once the three events are jointly considered.
@@
Fig. 3 shows, in blue, the two-dimensional $90\%$ credible region for $\mu_{B}$ and $\zeta$ for the $N_{\rm bosons}=1$ model, together with the corresponding one-dimensional posterior distributions obtained when we exclude S200114. We obtain a boson mass of $\mu_{B}=8.64^{+4.84}_{-7.17}\times 10^{-13}\,\rm{eV}$ and a fraction of $\zeta=0.28^{+0.43}_{-0.25}$ at the $90\%$ credible level. The value of $\mu_{B}$ is consistent with that obtained from the individual analysis of GW190521 in Bustillo et al. (2021). Notably, however, the uncertainty is significantly larger due to the fact that $\zeta$ is consistent with 0, in which case no constraints can be imposed on $\mu_{B}$.
@@
The solid-green line in Fig. 2 shows that the inclusion of S200114 does not change our qualitative conclusions. Quantitatively, however, it makes the $N=0$ and $N=1$ become equally preferred. The reason is that, while there now are two events favouring the PSM hypothesis, these point to inconsistent boson masses. On the one hand, this makes the $N_{\rm bosons}=1$ model having to decide whether GW190521 or S200114 is better PSM candidate – as both cannot be at the same time if only one boson mass is allowed – preventing a larger evidence in favour of such model. The corresponding posterior distributions for $\mu_{B}$ and $\zeta$ for the $N_{\rm bosons}=1$ case are shown in orange in Fig. 3. The most remarkable difference with respect to the case omitting GW190521 is that, as expected, the inferred boson mass peaks now at the value obtained from the individual analysis of S200114, as this event has a larger preference for the PSM model than GW190521.
@@
The right and left panels of Fig. 4 show respectively our posterior distributions for $\mu_{1,2}$ and $\zeta_{1,2}$ for the case $N_{\rm bosons}=2$ where we allow for the existence of two bosons. To give an idea of the “level of preference” for each model, we also show the posterior for $\zeta_{1}+\zeta_{2}$. Solid and dashed curves respectively correspond to analyses excluding and including S200114. We highlight two main aspects. First, the posteriors for $\mu_{1,2}$ are now clearly inconsistent with each other when S200114 is included in the analysis, each being respectively consistent with the mass values of S200114 and GW190521. Second, the evidence for the $N_{\rm bosons}=2$ model is however slightly smaller than for $N_{\rm bosons}=1$. This happens because the increase of the likelihood produced by the fact that each of the two events can now be fit through a different boson mass cannot overcome the increased Occam Penalty paid by the model due to two extra degrees of freedom. This result may, at first sight, seem in contradiction with the fact that the posterior for $\zeta_{1}+\zeta_{2}$ clearly peaks away from zero, where it achieves null values. The interpretation of such posteriors for model selection, however, needs to fold in the fact that these are heavily influenced by the corresponding prior, which peaks at $\zeta_{1}+\zeta_{2}=1$. In fact, a rough estimation of the Bayes Factor for the $N_{\rm bosons}=2$ v.s. $\zeta_{1}+\zeta_{2}=0$ (i.e. $N_{\rm bosons}=0$) model via the Savage-Dickey ratio yields values between $1.3$ and $1.7$ favouring $N_{\rm bosons}=0$, which is consistent with the values shown in Fig. 2 (green).
@@
Switching to a uniform distance prior has an tremendous impact in our analysis. As mentioned before, this is due to the fact that such prior does not penalise the intrinsic weakness of the sources for which PSM simulations exist to date. This change causes both GW190521 and S200114 to strongly favour the PSM hypothesis.For instance, even if S200114 is excluded, we obtain ${\cal{B}}^{n=1}_{n=0}\simeq 8$ when exploiting boson-mass coincidences and ${\cal{B}}^{n\neq 0}_{n=0}\simeq 3$ when ignoring them. If S200114 is included then the $N_{\rm bosons}=2$ model is preferred due to its boson mass being inconsistent with that of GW190521. In particular, we obtain ${\cal{B}}^{n=2}_{n=0}\simeq 50$ and ${\cal{B}}^{n\neq 0}_{n=0}\simeq 40$ and ${\cal{B}}^{n=2}_{n=1}\simeq 1.3$. Needless to say, these results do not imply by any means an actual detection of PSMs. These, however, underscore the potential impact that the production of more realistic simulations of louder numerical simulations of PSMs – which would not be penalised by a realistic distance prior – could have in future studies searching for these objects.
Figure 5: Accumulated evidence for Proca-star mergers as a function of the number of detected events. Left: $90\%$ credible intervals for the natural log Bayes Factor for the $N_{\rm bosons}=1$ (red) and $N_{\rm bosons}=2$ (green) bosons v.s. $N_{\rm bosons}=0$ as a function of the number of times that event sets like GW190521+GW190426+GW200220 are observed. The blue contour shows the evidence for non-zero v.s. zero Proca-star mergers when potential boson-mass coincidences across events are ignored. The magenta contour shows the log Bayes Factor for $N_{\rm bosons}=1$ v.s. $N_{\rm bosons}=2$. Right: Same as in the right panel, including the event S200114 in the event sets. Here, the magenta contour denotes the log Bayes factor for $N_{\rm bosons}=2$ v.s. $N_{\rm bosons}=1$.
@@ IV Detecting Proca-stars in mock data sets
We investigate the power of our method to extract evidence for a population of PSMs and to estimate the corresponding boson mass(es) as observations accumulate. To do so, we simulate the observation of $N_{\rm obs}$ instances of the event sets discussed previously. Each simulated observation retains the same relative evidences for the BBH and PSM models as the original events, but incorporates statistical fluctuations by modifying the boson-mass posterior distributions. Specifically, we generate perturbed posteriors $p^{*}(\mu_{B})$ by shifting the original posteriors $p(\mu_{B})$ by random offsets $\delta_{\mu_{B}}$ , sampled from the distribution $p(\mu_{B}-\mu^{50}_{B})$, where $\mu^{50}_{B}$ denotes the median of the original posterior. This approach introduces realistic noise while preserving the structure of the original inference.
@@
Figure 5 shows the $90\%$ credible intervals for the Bayesian evidence $p(\{d_{i}\}|n\equiv N)$ for the number of bosons $n=N_{\rm bosons}$ as a function of the number $N_{\rm obs}$ of repeated observations of a given set of events. In particular, the left panel ignores S200114 while the right one includes it in the event sets. We respectively represent in red and green the natural logarithm of the ratio of the evidences for $N_{\rm bosons}=1,2$ models v.s. the $N_{\rm bosons}=0$ model, together with the $N_{\rm bosons}=2$ v.s. $N_{\rm bosons}=1$ ratio in magenta. In addition, we show in blue the result for the $N_{\rm bosons}\neq 0$ v.s. $N_{\rm bosons}=0$, where the $N_{\rm bosons}\neq 0$ model ignores mass consistencies across events. The left panel shows results corresponding to $N_{\rm obs}$ instances of the 3-event sent ignoring S200114, while the right panel includes it.
The blue contour of the left panel clearly shows that, if mass-consistencies are ignored, the $N_{\rm bosons}\neq 0$ model is progressively discarded as the number of observations increases. On the contrary, the green and red contours show that if mass-consistencies are taken into account, evidence for the $N_{\rm bosons}\neq 0$ models are progressively favoured. In particular, conclusive evidence is obtained for the $N_{\rm bosons}=1$ model, with $\log{\cal B}^{n=1}_{n=0}>5$ at the $90\%$ level after $N_{\rm obs}\simeq 8$ observations of the event set. The magenta contour shows that the preference for the $N_{\rm bosons}=1$ over the $N_{\rm bosons}=2$ one progressively increases as events accumulate. However, given the properties of the events involved (very similar masses and only non very accurate mass estimates), obtaining conclusive preference for $N_{\rm bosons}=1$ would require a very large number of repeated observations. In fact, we have checked that the $N_{\rm bosons}=2$ model cannot be conclusively ruled out even if $N_{\rm obs}=50$.
@@
The progressive preference for the $N_{\rm bosons}=1$ model is accompanied by progressively tighter estimates of the boson mass, whose $90\%$ credible interval is shown in the left panel of Fig. 6 (blue). The green and yellow contours of the same panel denote the boson mass estimates obtained with the $N_{\rm bosons}=2$ model. As somewhat expected, such estimates have large uncertainties that always include the GW190521 median mass within the $90\%$ credible interval. This is due to the accuracy of the mass estimate being below the allowed minimum difference $\delta\mu_{B}$ between different masses. In other words, one cannot distinguish between a single boson model and a model with two bosons with extremely similar masses that differ by less than our maximum allowed resolution $\delta\mu_{B}$.
@@
The right panels of Figs. 5 and 6 shows the same results discussed above but for the case where S200114 is included in the analysis. The only exception is that the magenta contour in the right panel of Fig. 5 now corresponds to the ratio of the evidences between the $N_{\rm bosons}=2$ and $N_{\rm bosons}=1$ models. The inclusion of S200114 and the corresponding copies in the event sets has several implications. First, conclusive evidence for $N_{\rm bosons}\neq 0$ can be obtained now after $\simeq 8$ repeated observations even if mass-coincidences are ignored due to the higher preference of S200114 for the PSM model. Second, such evidence is obtained after only $\simeq 4$ repeated observations if mass-coincidences are included using the $N_{\rm bosons}=2$ model. Finally, conclusive evidence for the $N_{\rm bosons}=2$ v.s. the $N_{\rm bosons}=1$ model is obtained after $\simeq 9$ repeated observations.
@@
In terms of boson-mass estimations, the right panel of Fig. 6 shows that the two masses considered in the $N_{\rm bosons}=2$ model progressively converge to the median values of S200114 and GW190521, with the former being more accurately estimated right from the start as a consequence of the larger preference for the PSM model. Additionally, we note that, given its restriction to a single boson, the $N_{\rm bosons}=1$ model clearly discards GW190521 and its copies as a PSMs, ignoring the corresponding boson mass estimates; and opting instead to classify S200114 as a PSM, as somewhat expected.

### v2

@@ II.3 A boson-counting experiment:
We now generalise the above formalism to a case where we allow for $n$ Proca-star families, each characterised by a value of the boson mass $\mu_{B}^{j}$ with $j\in[1,n]$. In this case, the joint likelihood for the $n$ boson masses $\mu_{B}^{j}$ and the corresponding fractions $\zeta_{j}$ of the $N$ analysed events belonging to each family within the observed data is given by
@@
$=\prod_{i=1}^{N}\left[\sum_{j=1}^{n}p(d_{i}|\text{PSM}(\mu_{B}^{j}))\zeta_{j}+p(d_{i}|\text{BBH})\quantity(1-\sum_{j=1}^{n}\zeta_{j})\right]$ $\propto\prod_{i=1}^{N}\quantity[\sum_{j=1}^{n}{\cal{B}}^{\text{PSM}}_{\text{BBH},i}(\mu_{B}^{j})\zeta_{j}+\quantity(1-\sum_{j=1}^{n}\zeta_{j})].$
@@ V D V D V D V D
Finally, the Bayesian evidence for a model allowing for a number of bosons $n$ is just given by
[…]
where $i\in[1,N]$.
With this, we can construct the posterior distribution on the number of bosons $n$ as
@@
where we will always consider an uniform prior probability $\pi(n)$ on $n$. Finally, we note that the likelihood for the case $n=0$, where all events are considered to be BBHs, is simply given by setting $\zeta_{i}=0$ in Eq. (5), which is equivalent to setting $\zeta=0$ in Eq. (1).
@@ II.4 Parameter sampling
We impose uniform priors on the boson masses with $\mu_{B}^{i}\in[2,15]\times 10^{-13}$ eV. This range safely contains the full posterior distributions for $\mu_{B}$ individually inferred for the four events we study [^1] We also note that setting $\mu_{B}=0$ reduces the Proca equation to Maxwell’s equations, corresponding to the massless photon case. In this limit the field supports only long-range radiative solutions, which prevents the existence of localized, gravitationally bound configurations that could form compact objects.. In order to label the different bosons and avoid double-counting, we choose to impose $\mu^{i+1}_{B}\leq\mu^{i}_{B}$. Importantly, for computational-cost reasons, we do not sample over a continuous range of values $\mu^{i}_{B}$. Instead, we sample over a discrete grid with step $\delta\mu_{B}=1\times 10^{-15}$ eV. Physically, $\delta\mu_{B}$ can be interpreted as the maximum resolution in the boson mass we allow for, which is much smaller than the standard deviation of the individual posteriors for each event that we take as input. We also set a flat prior on the respective fractions $\zeta_{i}\in[0,1]$, imposing $\sum_{i}\zeta_{i}\leq 1$. We sample the parameter space using the nested sampler Dynesty Speagle (2020) with 4096 live points.
@@ II.5 Inputs for our analysis
We take as input data the Bayesian evidences $p(d_{i}|{\rm BBH})$ and $p(d_{i}|{\rm PSM})$ for the BBH and PSM models for the four events we consider reported in Calderón Bustillo et al. (2023), together with the corresponding posterior distributions $p(\mu_{B}|d_{i})$ for the boson mass obtained under the PSM model. These are respectively shown in Fig. 1 and Table 1. In the mentioned study, the authors performed Bayesian parameter inference and model selection on the events we consider, both under the BBH and PSM scenarios. For the BBH case, the data was analyzed with the state-of-the art waveform model for BBHs NRSur7dq4 Varma et al. (2019). This model includes waveform multipoles up to $\ell=4$. The model is directly calibrated to numerical simulations of generically spinning BBHs Boyle et al. (2019) with mass ratios $q\in[1,4]$ and spin-magnitudes $a_{i}\in[0,0.8]$ but can be extrapolated to $q=6$ and extremal spins. For the PSM case, the authors employed a family of 769 numerical simulations of unequal-mass head-on mergers of Proca stars, with the corresponding field frequencies ranging in $\omega/\mu_{B}\in[0.80,0.93]$ and imposing a flat prior on these parameters. In these configurations, the stars were released at rest at a separation distance of 20M20\,M, expressed in geometric units. For a visualization of the GW detector data for the four events we study, together with the corresponding best-fitting BBH and PSM waveforms, we refer the reader to Figs. 1-4 in Calderón Bustillo et al. (2023).
@@ III Results on real data
Figure 2: Posterior probability for the number $N_{\rm bosons}$ of ultra-light bosons given the Advanced LIGO - Virgo events we consider. Solid (dashed) lines ignore (include) the event S200114 in the analysis. Blue and green lines respectively make use of distance priors uniform in co-moving volume and on luminosity distance. Note that, in order to alleaviate notation, in the main text we denote the variable $N_{\rm bosons}$ simply by $n$.
@@
Figure 3: Posterior distributions for the boson-mass and Proca-star fraction obtained when we only allow for a single ultralight boson. Blue and orange curves and contours denote results obtained by respectively including and excluding and including the S200114 trigger. The results are obtained assuming a distance prior uniform in co-moving volume. The posterior distribution for the boson mass has non-zero support for all the prior range because the posterior distribution for $\zeta$ is non-null at $\zeta=0$, in which case all values of $\mu_{B}$ are allowed.
@@
Figure 4: Posterior distributions for boson-mass (right) and Proca-star fractions (left) for the 2-boson model. Solid and dashed curves denote results respectively excluding and including the S200114 trigger. Blue and orange curves respectively denote the value of the primary and secondary boson-mass $\mu_{B}^{1,2}$ and the corresponding fraction $\zeta_{1,2}$. Black curves denote the posterior distribution for $\zeta_{1}+\zeta_{2}$.
@@
Fig. 2 shows the posterior probability for the number of ultralight bosons $n$ given our different data sets and priors. Blue and green curves respectively exclude and include the trigger S200114. We start by discussing results obtained under the regular distance prior. In this situation, the $n=0$ model is preferred with a Bayes factor ${\cal{B}}^{n=0}_{n=1}\simeq 1.6$ with respect to the $n=1$ model. This weak preference for $N=0$ is expected, as Table 1 shows that only one of the analyzed events has a slight preference for the PSM scenario while the other two reject such hypothesis rather strongly. In other words, the slight preference of GW190521 for the PSM hypothesis is consistent with a statistical fluctuation once the three events are jointly considered.
@@
Fig. 3 shows, in blue, the two-dimensional $90\%$ credible region for $\mu_{B}$ and $\zeta$ under the $n=1$ model obtained when we exclude S200114, together with the corresponding one-dimensional posterior distributions. We obtain a boson mass $\mu_{B}=8.64^{+4.84}_{-7.17}\times 10^{-13}\,\rm{eV}$ and a fraction of $\zeta=0.28^{+0.43}_{-0.25}$ at the $90\%$ credible level. The value of $\mu_{B}$ is consistent with that obtained from the individual analysis of GW190521 in Bustillo et al. (2021). Notably, however, the uncertainty is significantly larger due to the fact that $\zeta$ is consistent with 0, in which case no constraints can be imposed on $\mu_{B}$.
@@
The solid-green line in Fig. 2 shows that the inclusion of S200114 does not change our qualitative conclusions. Quantitatively, however, it makes the $n=0$ and $n=1$ become equally preferred. The reason is that, while there are now two events favouring the PSM hypothesis, these point to inconsistent boson masses. This makes the $n=1$ model having to decide whether GW190521 or S200114 is a better PSM candidate – as both cannot be at the same time if only one boson mass is allowed – preventing a larger evidence in favour of such model. The corresponding posterior distributions for $\mu_{B}$ and $\zeta$ for the $n=1$ case are shown in orange in Fig. 3. The most remarkable difference with respect to the case omitting GW190521 is that, as expected, the inferred boson mass peaks now at the value obtained in the individual analysis of S200114, as this event has a larger preference for the PSM model than GW190521.
@@
The right and left panels of Fig. 4 show respectively our posterior distributions for $\mu^{1,2}_{B}$ and $\zeta_{1,2}$ for the case $n=2$ where we allow for the existence of two bosons. To give an idea of the “level of preference” for this model v.s. the $n=0$ one, we also show the posterior for $\zeta_{1}+\zeta_{2}$. Solid and dashed curves respectively correspond to analyses excluding and including S200114. We highlight two main aspects. First, the posteriors for $\mu^{1,2}_{B}$ are clearly inconsistent with each other when S200114 is included in the analysis, each being respectively consistent with the mass values of S200114 and GW190521. Second, as shown in Fig. 2 the evidence for the $n=2$ model is however slightly smaller than for $n=1$. This happens because the increase of the likelihood produced by the fact that each of the two events can now be fit through a different boson mass cannot overcome the increased Occam Penalty paid by the two extra degrees of freedom of this model. This result may, at first sight, seem in contradiction with the fact that the posterior for $\zeta_{1}+\zeta_{2}$ clearly peaks away from zero, where it achieves null values. The interpretation of such posteriors for model selection, however, needs to fold in the fact that these are heavily influenced by the corresponding prior, which peaks at $\zeta_{1}+\zeta_{2}=1$. In fact, a rough estimation of the Bayes Factor for the $n=2$ v.s. $\zeta_{1}+\zeta_{2}=0$ (i.e. $n=0$) model via the Savage-Dickey ratio yields values between $1.3$ and $1.7$ favouring $n=0$. This is consistent with the values shown in Fig. 2 (green).
@@
We now study the impact of the intrinsic weakness of head-on PSM mergers in the above results by considering a prior uniform in distance that removes such effect from our analysis. While this prior and the corresponding conclusions in terms of parameter inference not physically meaningful, it allows us to understand how results under our regular prior are impacted by the intrinsic weakness of head-on PSMs. We indeed see that this has a tremendous impact in our analysis. Switching to this prior causes both GW190521 and S200114 to strongly favour the PSM hypothesis. As a consequence, even if S200114 is excluded from the analysis, we obtain ${\cal{B}}^{n=1}_{n=0}\simeq 8$ when exploiting boson-mass coincidences and ${\cal{B}}^{n\neq 0}_{n=0}\simeq 3$ when ignoring them. If S200114 is included then the $n=2$ model is preferred due to its boson mass being inconsistent with that of GW190521. In particular, we obtain ${\cal{B}}^{n=2}_{n=0}\simeq 50$ and ${\cal{B}}^{n\neq 0}_{n=0}\simeq 40$ and ${\cal{B}}^{n=2}_{n=1}\simeq 1.3$. Needless to say, these results do not imply by any means an actual detection of PSMs. These, however, underscore the potential impact that the production of more realistic simulations of louder PSMs configurations – which would not be penalised by a realistic distance prior – could have in future studies searching for these objects.
Figure 5: Accumulated evidence for Proca-star mergers as a function of the number of observations. Left: $90\%$ credible intervals for the natural log Bayes Factor for the $N_{\rm bosons}=1$ (red) and $N_{\rm bosons}=2$ (green) bosons v.s. $N_{\rm bosons}=0$ as a function of the number of times that event sets like GW190521+GW190426+GW200220 are observed. The blue contour shows the evidence for non-zero v.s. zero Proca-star mergers when potential boson-mass coincidences across events are ignored. The magenta contour shows the log Bayes Factor for $N_{\rm bosons}=1$ v.s. $N_{\rm bosons}=2$. Right: Same as in the right panel, including the event S200114 in the event sets. Here, the magenta contour denotes the log Bayes factor for $N_{\rm bosons}=2$ v.s. $N_{\rm bosons}=1$.
@@ IV Detecting Proca-stars in mock data sets
In this section, we investigate the power of our framework to extract evidence for a population of PSMs and to estimate the corresponding boson mass(es) as observations accumulate. In particular, we want to consider events similar to the ones at hand, which satisfy that a) no event can individually be identified as a PSMs in a conclusive way b) several of these events actually point to similar boson masses. To do so, we simulate the observation of $N_{\rm obs}$ instances of the two sets of real events we have analyzed in this work, respectively ignoring and including simulated events similar to S200114f. Each simulated observation will retain the same relative evidences for the BBH and PSM models as the events we have analyzed, but incorporates statistical fluctuations in the corresponding posterior distribution of the boson-mass. Specifically, we generate such perturbed posteriors $p^{*}(\mu_{B})$ by shifting the original posteriors $p(\mu_{B})$ by random offsets $\delta_{\mu_{B}}$, sampled from the distribution $p(\mu_{B}-\mu^{50}_{B})$, where $\mu^{50}_{B}$ denotes the median of the original posterior. Notably, this approach also allows us to forecast “how many instances of the studied events we would need to observe in order to claim a Proca-star detection” respectively using frameworks that exploit and ignore boson-mass coincidences.
@@
Figure 5 shows the $90\%$ credible intervals for the Bayesian evidence $p(\{d_{i}\}|n)$ for the number of bosons $n$ as a function of the number $N_{\rm obs}$ of repeated observations of a given set of events. In particular, the left panel ignores simulated events similar to S200114 while the right one includes them in the event sets. We respectively represent in red and green the natural logarithm of the ratio of the evidences for $n=1,2$ models v.s. the $n=0$ model, together with the $n=2$ v.s. $n=1$ ratio in magenta. In addition, we show in blue the result for the $n\neq 0$ v.s. $n=0$, where the $n\neq 0$ model ignores mass consistencies across events. The left panel shows results corresponding to $N_{\rm obs}$ instances of the 3-event sent ignoring S200114, while the right panel includes it.
[…]
@@
The progressive preference for the $n=1$ model is accompanied by progressively tighter estimates of the boson mass, whose $90\%$ credible interval is shown in the left panel of Fig. 6 (blue). The green and yellow contours of the same panel denote the boson mass estimates obtained with the $n=2$ model. As somewhat expected, such estimates have large uncertainties that always include the GW190521 median mass within the $90\%$ credible interval. This is due to the accuracy of the mass estimate being below the allowed minimum difference $\delta\mu_{B}$ between different masses. In other words, one cannot distinguish between a single boson model and a model with two bosons with extremely similar masses that differ by less than our maximum allowed resolution $\delta\mu_{B}$.
@@
The right panels of Figs. 5 and 6 shows the same results discussed above but for the case where S200114 is included in the analysis. The only exception is that the magenta contour in the right panel of Fig. 5 now corresponds to the ratio of the evidences between the $n=2$ and $n=1$ models. The inclusion of events similar to S200114 in the event sets has several implications. First, conclusive evidence for $n\neq 0$ can be obtained now after $\simeq 8$ repeated observations even if mass-coincidences are ignored due to the higher preference of S200114 for the PSM model. Second, such evidence is obtained after only $\simeq 4$ repeated observations if mass-coincidences are included using the $n=2$ model. Finally, conclusive evidence for the $n=2$ v.s. the $n=1$ model is obtained after $\simeq 9$ repeated observations.
@@
In terms of boson-mass estimations, the right panel of Fig. 6 shows that the two masses considered in the $n=2$ model progressively converge to the median values of S200114 and GW190521, with the former being more accurately estimated right from the start as a consequence of its larger preference for the PSM model. Additionally, we note that, given its restriction to a single boson, the $n=1$ model clearly discards GW190521 and similar simulated events as a PSMs, ignoring the corresponding boson mass estimates; opting instead to classify S200114 as a PSM, as somewhat expected.